# Attention as a Hypernetwork

### v3

@@ 7 Discussion
We have proposed a novel decomposition of multi-head attention revealing that it can be interpreted as a hypernetwork that composes value networks specific to each key-query pair. Consistent with this perspective, we find empirically that the attention scores over the heads for each key-query pair form a functionally structured space, identifying reusable subfunctions in two abstract reasoning tasks. Furthermore, modifying multi-head attention to strengthen the hypernetwork mechanism improves compositional generalization on these tasks.
Adopting the hypernetwork perspective more broadly, multi-head attention can be interpreted as a particular choice for the level of granularity at which the hypernetwork parameterizes the value network: The value networks are key-query specific and are followed by a pooling step that sums the key-query specific outputs over the key index. In principle, other levels of granularity are possible. For instance the hypernetwork could parameterize a query specific value network which subsumes the aggregation over the key index. In the extreme case, it could directly parameterize the full sequence operation potentially facilitating the re-use of sequence-level operations. This also offers an interesting connection to attention-based graph neural networks [44, 45]. Classically, a non-causal single layer transformer is seen as a fully connected graph where each token corresponds to a node and aggregation is done via attention [46]. Our derivation suggests an alternative interpretation where the message function is a hypernetwork that subsumes the attention weights while the aggregation becomes the sum operator. Given the crucial role that aggregation plays in graph neural networks [47, 48], a natural question is whether different pooling operators should be considered more generally for multi-head attention.
@@
Acknowledgements.
We would like thank Angelika Steger as well as Alexander Meulemans, Johannes von Oswald, Maciej Wołczyk and Owen He for fruitful discussions throughout the development of the project. This research was supported by an Ambizione grant (PZ00P3_186027) from the Swiss National Science Foundation and an ETH Research Grant (ETH-23 21-1).
@@ A.1 Model ablations
Compared to linear attention, HYLA introduces three modifications: (1) normalization of the attention scores across the head indices using RMSNorm, (2) applying the attention-weighted sum operation also to the output projection and (3) inserting a nonlinearity in the value network (c.f. the pseudocode for HYLA shown in Figure 1). To delineate the contributions of these modifications, we run an ablation study on the fuzzy logic task and sraven shown in Table A1. We find that while simply applying RMSHead to linear attention (Linear Attention + RMSHead) and also the attention-weighted sum operation without nonlinearity (HYLA -- nonlinearity) both by themselves can slightly boost performance, the full HYLA model generally performs best.
@@
We perform an additional ablation where we replace the RMSHead normalization of HYLA with the standard softmax. As shown in Table A1 this model performs poorly. This might be partially explained by the fact that with the softmax the linear combination of weights produced by the hypernetwork is no longer guaranteed to maintain the scale of a standard neural network initialization corresponding to the value network.
@@
Figure A2: Training coefficient of determination ($R^{2}$) for varying sequence length and fraction of held out tasks when in-context learning of fuzzy logic functions.
@@
Figure A4: Performance on fuzzy logic functions containing terms not encountered during training. As expected all methods fail in this setting. Same setup as in Figure 2B.
@@ B.1 Number of possible tasks and instances
Figure A7: A Number of possible sraven tasks as a function the number of features ($M$). B Number of possible sraven problem instances as a function of the number of features ($M$) and the number of feature values ($K$).
In the general case, the number of possible tasks in sraven is a product of the number of possible rule combinations $\binom{R+M-1}{M}$ and the number of possible unique permutations $((M!)^{G-1}-M)$ where $G$ is the number of columns of the grid (in Raven-based tasks typically $G=3$). Figure A7A shows the number of possible sraven tasks and Figure A7B shows the number of possible sraven problem instances. Note specifically that the latter is orders of magnitudes larger than the number of training examples in our experiments for $M=4$, $K=8$.
@@ B.2 Ambiguity
Number of features ($M$) Number of feature values ($K$) Estimated fraction of ambiguous examples
@@
In principle, a problem instance generated given a set of rules and permutations might have an ambiguous answer. That is there is at least one other set of rules and permutations that fit all context panels but prescribe a different final query panel. One possibility is to check for such examples when generating task instances and reject ambiguous ones. This process is prohibitively expensive as it requires a search over the exponentially many possible permutations. We therefore resort to a Monte-Carlo estimation of the fraction of ambiguous instances. Table A2 shows the results for various parameterizations of the sraven task. In the setting we consider here, ($M=4,K=8$), the probabilities are negligible and we do not take measures to filter out ambiguous examples given the computational overhead it produces in the data generation process.
@@ B.3 Rules
    constant: Each row consists of a random but fixed integer in the range $1$ to $K$.
@@
    progression (+1): The first element of each row is sampled uniformly at random and incremented by 1 modulo $K$ for each successive column.
@@
    progression (+2): The first element of each row is sampled uniformly at random and incremented by 2 modulo $K$ for each successive column.
@@
    progression (-1): The first element of each row is sampled uniformly at random and decremented by 1 modulo $K$ for each successive column.
@@
    progression (-2): The first element of each row is sampled uniformly at random and decremented by 2 modulo $K$ for each successive column.
@@
    addition: Two elements are sampled uniformly at random for each row and added modulo $K$ to obtain the last column.
@@
    subtraction: Two elements are sampled uniformly at random for each row and subtracted modulo $K$ to obtain the last column.
@@ Appendix C Experimental details
MultiHeadAttention is either linear attention, softmax attention or hypernetwork linear attention (HYLA) and uses T5-style relative positional embeddings [33]. For the FeedForward layer, we employ a standard single-hidden layer MLP with GeLU nonlinearity applied to each position in the sequence independently. For our language modeling experiments we use rotary positional encodings [49].
@@
For the fuzzy logic task we directly use the concatenated input examples as tokens. Similarly for sraven we use one-hot encodings of the integer inputs as the input tokens. In both cases the input tokens are mapped into the embedding space of the transformer using a fully connected dense layer and another fully connect dense layer is used to produce the output logits. For the language modeling experiment we use the default sentencepiece tokenizer [50] of NanoDo [51] with a vocabulary size of $32000$ and use the transposed embedding matrix to produce the logits.
@@
We use the AdamW optimizer [52] with linear warmup starting from a learning rate of 00 followed by a cosine decay learning rate scheduler [53] that decays the learning rate to $0.1$ times the base learning rate. We exempt biases and LayerNorm parameters from weight decay. In the fuzzy logic task we use the mean-squared error on the logits of the query token whereas in sraven we use the softmax cross-entropy loss on the logits of all query tokens. The language modeling task uses an autoregressive softmax cross-entropy loss with causal masking applied to the attention mechanism.
@@ D.2 Software and libraries
For the results obtained in this paper we built on free and open-source software. We implemented our experiments in Python using JAX [54, Apache License 2.0], Flax [55, Apache License 2.0], NanoDo [51, Apache License 2.0] and the Deepmind Jax Ecosystem [56, Apache License 2.0]. We utilized WandB [57, MIT license] to monitor the progress and results of experiments, and Plotly [58, MIT license] for generating the plots.

### v4

@@ 7 Discussion
We have proposed a novel decomposition of multi-head attention, revealing that it can be interpreted as a hypernetwork that composes value networks specific to each key-query pair. Consistent with this perspective, we find empirically that the attention scores over the heads for each key-query pair form a functionally structured space, identifying reusable subfunctions in two abstract reasoning tasks. Furthermore, modifying multi-head attention to strengthen the hypernetwork mechanism improves compositional generalization on these tasks.
Adopting the hypernetwork perspective more broadly, multi-head attention can be interpreted as a particular choice for the level of granularity at which the hypernetwork parameterizes the value network: The value networks are key-query specific and are followed by a pooling step that sums the key-query specific outputs over the key index. In principle, other levels of granularity are possible. For instance, the hypernetwork could parameterize a query specific value network which subsumes the aggregation over the key index. In the extreme case, it could directly parameterize the full sequence operation potentially facilitating the re-use of sequence-level operations. This also offers an interesting connection to attention-based graph neural networks (Veličković et al., 2018; Shirzad et al., 2023). Classically, a non-causal single layer transformer is seen as a fully connected graph where each token corresponds to a node and aggregation is done via attention (Battaglia et al., 2018). Our derivation suggests an alternative interpretation where the message function is a hypernetwork that subsumes the attention weights while the aggregation becomes the sum operator. Given the crucial role that aggregation plays in graph neural networks (Rosenbluth et al., 2023; Dudzik et al., 2023), a natural question is whether different pooling operators should be considered more generally for multi-head attention.
@@
Acknowledgements
We would like to thank Angelika Steger as well as Alexander Meulemans, Johannes von Oswald, Maciej Wołczyk and Owen He for fruitful discussions throughout the development of the project. This research was supported by an Ambizione grant (PZ00P3_186027) from the Swiss National Science Foundation and an ETH Research Grant (ETH-23 21-1).
@@ A.1 Model ablations
Compared to linear attention, HYLA introduces three modifications: (1) normalization of the attention scores across the head indices using RMSHead, (2) applying the attention-weighted sum operation also to the output projection and (3) inserting a nonlinearity in the value network (c.f. the pseudocode for HYLA shown in Figure 1). To delineate the contributions of these modifications, we run an ablation study on the fuzzy logic task and sraven shown in Table A1. We find that while simply applying RMSHead to linear attention (Linear Attention + RMSHead) and also the attention-weighted sum operation without nonlinearity (HYLA -- nonlinearity) both by themselves can slightly boost performance, the full HYLA model generally performs best.
@@
We perform an additional ablation where we replace the RMSHead normalization of HYLA with the standard softmax that normalizes across keys but not across heads. As shown in Table A1 this model performs poorly.
Replacing feedforward layers with mixture of experts.
Sparsely gated mixture of experts (MoE) have been proposed as a replacement for the feedforward layer in the transformer block (Shazeer et al., 2017). In this model, each token is routed to a sparse subset of $k$ out of $E$ experts
$f_{\mathrm{MoE}}(\bm{x})=\sum_{i=1}^{E}\mathrm{Softmax}(\mathrm{Top\_k}(\bm{W}^{\text{gate}}\bm{x}))\cdot\bm{W}^{\mathrm{up}}_{i}\phi(\bm{W}^{\mathrm{down}}_{i}\bm{x}).$
Intuitively, this could allow individual experts to specialize on specific subtasks that can be composed by the router. However, it is not well understood if MoEs improve compositional generalization of transformers in practice. Here, we perform additional experiments where we replace the feedforward layers with MoEs on the sraven and the fuzzy logic task reported in Table A1. While we find moderate general improvements, these improvements seem to be similar to other increases of model capacity such as increasing the width/depth as discussed in Section 4.4.
@@
Softmax attention + MoE $63.83\pm 6.07$ $57.68\pm 1.31$
Linear attention + MoE $60.09\pm 6.16$ $64.34\pm 4.3$
HYLA + MoE $82.37\pm 6.55$ $66.94\pm 3.94$
@@
Figure A3: Fuzzy logic task performance metrics. Various task performance metrics on the fuzzy logic task for varying number of in-context examples and fraction of tasks held-out during training. A Coefficient of determination ($R^{2}$) on tasks consisting of unseen combinations of seen terms (same as Figure 2C) B Coefficient of determination ($R^{2}$) on training tasks. C Coefficient of determination ($R^{2}$) on tasks containing terms not encountered during training. As expected all methods fail in this setting. D Training mean squared error loss.
Figure A4: Training loss when scaling data and model size on sraven. A Training loss as a function of the number of training problem instances for different widths (scaling embedding and key-query-value dimensions). B Same as A but increasing depth instead of width.
Figure A5: Varying the fraction of OOD tasks on sraven. OOD accuracy on sraven as varying fractions of tasks specified by their rule combinations are held-out during training.
@@
Figure A8: Detailed attention scores on sraven. A Random sample of the attention scores / latent codes for the final key-query pair on OOD sraven tasks in the final layer of a transformer using softmax attention, linear attention or HYLA. B Average attention scores / latent codes per rule. C Pair-wise cosine similarities between all average latent codes per rule. Data from the same models as shown in Figure A6)
@@ B.1 Number of possible tasks and instances
Figure A9: A Number of possible sraven tasks as a function of the number of features ($K$). B Number of possible sraven problem instances as a function of the number of features ($K$) and the number of feature values ($F$).
In the general case, the number of possible tasks in sraven is a product of the number of possible rule combinations $\binom{R+K-1}{K}$ and the number of possible unique permutations $((K!)^{G-1}-K)$ where $G$ is the number of columns of the grid (in Raven-based tasks typically $G=3$). Figure A9A shows the number of possible sraven tasks and Figure A9B shows the number of possible sraven problem instances. Note specifically that the latter is orders of magnitudes larger than the number of training examples in our experiments for $K=4$, $F=8$.
@@ B.2 Ambiguity
Number of features ($K$) Number of feature values ($F$) Estimated fraction of ambiguous examples
@@
In principle, a problem instance generated given a set of rules and permutations might have an ambiguous answer. That is, there is at least one other set of rules and permutations that fit all context panels but prescribe a different final query panel. One possibility is to check for such examples when generating task instances and reject ambiguous ones. This process is prohibitively expensive, as it requires a search over the exponentially many possible permutations. We therefore resort to a Monte-Carlo estimation of the fraction of ambiguous instances. Table A2 shows the results for various parameterizations of the sraven task. In the setting we consider here, ($K=4,F=8$), the probabilities are negligible, and we do not take measures to filter out ambiguous examples given the computational overhead it produces in the data generation process.
@@ B.3 Rules
    constant: Each row consists of a random but fixed integer in the range $1$ to $F$.
@@
    progression (+1): The first element of each row is sampled uniformly at random and incremented by 1 modulo $F$ for each successive column.
@@
    progression (+2): The first element of each row is sampled uniformly at random and incremented by 2 modulo $F$ for each successive column.
@@
    progression (-1): The first element of each row is sampled uniformly at random and decremented by 1 modulo $F$ for each successive column.
@@
    progression (-2): The first element of each row is sampled uniformly at random and decremented by 2 modulo $F$ for each successive column.
@@
    addition: Two elements are sampled uniformly at random for each row and added modulo $F$ to obtain the last column.
@@
    subtraction: Two elements are sampled uniformly at random for each row and subtracted modulo $F$ to obtain the last column.
@@ Appendix C Experimental details
MultiHeadAttention is either linear attention, softmax attention or hypernetwork linear attention (HYLA) and uses T5-style relative positional embeddings (Raffel et al., 2020). For the FeedForward layer, we employ a standard single-hidden layer MLP with GeLU nonlinearity applied to each position in the sequence independently. For our language modeling experiments we use rotary positional encodings (Su et al., 2024).
@@
For the fuzzy logic task we directly use the concatenated input examples as tokens. Similarly for sraven we use one-hot encodings of the integer inputs as the input tokens. In both cases the input tokens are mapped into the embedding space of the transformer using a fully connected dense layer and another fully connect dense layer is used to produce the output logits. For the language modeling experiment we use the default sentencepiece tokenizer (Kudo & Richardson, 2018) of NanoDo (Liu et al., 2024) with a vocabulary size of $32000$ and use the transposed embedding matrix to produce the logits.
@@
We use the AdamW optimizer (Loshchilov & Hutter, 2019) with linear warmup starting from a learning rate of 00 followed by a cosine decay learning rate scheduler (Loshchilov & Hutter, 2017) that decays the learning rate to $0.1$ times the base learning rate. We exempt biases and LayerNorm parameters from weight decay. In the fuzzy logic task we use the mean-squared error on the logits of the query token whereas in sraven we use the softmax cross-entropy loss on the logits of all query tokens. The language modeling task uses an autoregressive softmax cross-entropy loss with causal masking applied to the attention mechanism as well as mixed-precision training.
@@ D.2 Software and libraries
For the results obtained in this paper we built on free and open-source software. We implemented our experiments in Python using JAX (Bradbury et al., 2018, Apache License 2.0), Flax (Heek et al., 2023, Apache License 2.0), NanoDo (Liu et al., 2024, Apache License 2.0) and the Deepmind Jax Ecosystem (Babuschkin et al., 2020, Apache License 2.0). We utilized WandB (Biewald, 2020, MIT license) to monitor the progress and results of experiments, and Plotly (Inc, 2015, MIT license) for generating the plots.